# A single-loop SPIDER-type stochastic subgradient method for expectation-constrained nonconvex nonsmooth optimization

### v1

@@ 4.1 Datasets and tested problems
(47) $\min_{{\mathbf{x}}}\Phi({\mathbf{x}})+\lambda\|\varphi({\mathbf{x}})\|_{1},\,\,\text{{s.t.}}\,\,\Psi_{0}({\mathbf{x}})\leq\kappa_{2},$
where $\varphi:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$, $\Psi_{0}:\mathbb{R}^{d}\mapsto\mathbb{R}$, $\Phi$ is the same as that in Eq. 4, $\lambda>0$, $\kappa_{2}>0$ is a slackness parameter. The functions $\Psi_{0}$ and $\varphi$ are defined as follows:
@@
$\small{[\varphi({\mathbf{x}})]_{i}=\begin{cases}2\left|x_{i}\right|,&0\leq\left|x_{i}\right|\leq 1,\\ -x_{i}^{2}+4\left|x_{i}\right|+1,&1<\left|x_{i}\right|\leq 2,\\ 3,&\left|x_{i}\right|>2,\end{cases}\quad\text{ for }i=1,2,\ldots,d}$
where $\varphi$ promotes sparsity in ${\mathbf{x}}$. Here, $\Psi_{0}$ is a continuous relaxation of the demographic parity $\left|\frac{1}{n_{\mathrm{p}}}\sum_{i=1}^{n_{\mathrm{p}}}\mathbb{I}\left(\mathbf{x}^{\top}\mathbf{a}_{i}^{\mathrm{p}}\geq 0\right)-\frac{1}{n_{\mathrm{u}}}\sum_{i=1}^{n_{\mathrm{u}}}\mathbb{I}\left(\mathbf{x}^{\top}\mathbf{a}_{i}^{\mathrm{u}}\geq 0\right)\right|$. For this problem, we have $\rho_{f}=\rho_{g}=\max\left\{2\lambda,\frac{1}{n_{\mathrm{p}}}\sum_{i=1}^{4n_{\mathrm{p}}}\|{\mathbf{a}}_{i}^{\mathrm{p}}\|^{2}+\frac{1}{n_{\mathrm{u}}}\sum_{i=1}^{4n_{\mathrm{u}}}\|{\mathbf{a}}_{i}^{\mathrm{u}}\|^{2}\right\}.$ In the experiments, we set $\lambda=0.02$, $\kappa_{2}=0.02$, and ${\mathbf{x}}^{(0)}=\mathbf{0}$ for all the compared methods.
@@ 4.2 Implementation details
For our method, we use $\beta=10$ and $\nu=10^{-5}$ as the default setting. We implement both deterministic and stochastic versions of our method (denoted as 3S-Econ-D and 3S-Econ-S, respectively). For 3S-Econ-D, we set $q=1$ and $S_{1}=S_{2}=n$ (i.e., taking the whole data) and select $\alpha_{k}=10^{-2}$ as our default setting. For 3S-Econ-S, we set $S_{1}=n$ and $S_{2}=q=\lceil\sqrt{n}\rceil$ and select $\alpha_{k}=\frac{1}{100\sqrt{k/q}}$ as our default setting [^2] To determine the default settings, we test 3S-Econ-D and 3S-Econ-S on Problem 1 using the a9a dataset, with varying choices of $\beta$ and $\alpha_{k}$. Specifically, for all $k\geq 0$, we choose $\beta\in\{1,10,100,1000\}$ and $\alpha_{k}\in\{0.001,0.01,\frac{1}{100\sqrt{k/q}},\frac{1}{1000\sqrt{k/q}}\}$. Through our experiments, we observe that 3S-Econ-D achieves the best performance with $\beta=10$ and $\alpha_{k}=0.01$. Similarly, for 3S-Econ-S, $\beta=10$ and $\alpha_{k}=\frac{1}{100\sqrt{k/q}}$ yield the best performance. Based on these observations, we fix these parameters as the default settings and apply them across all datasets and tested problems. .
@@
We adopt the tuning techniques in [9, 30] for the two distinct double-loop IPP methods and the techniques in [18] for the SSG. A strongly convex constrained subproblem is solved approximately in each outer iteration of the IPPs. For this subproblem, we employ the SSG from [18] and the ConEx in [9] as inner solvers. We refer to the two implemented IPPs as IPP-SSG and IPP-ConEx, respectively. The stepsize $\alpha_{k}$ for IPP-SSG is selected from $\{2\times 10^{-4},5\times 10^{-4},10^{-3},5\times 10^{-3}\}$. Following the notation in [30], we set $\theta_{t}=\frac{t}{t+1},\eta_{t}=c_{1}(t+1)$, and $\tau_{t}=\frac{c_{2}}{t+1}$ for IPP-ConEx, where $c_{1}$ is chosen from $\{20,50,100,200\}$ and $c_{2}$ from $\{0.002,0.005,0.01,0.02\}$. The optimal parameter set for these methods is selected based on the smallest objective value after 5000 iterations.
@@
We report the objective function value (FV), i.e., $f({\mathbf{x}})$, the constraint violation (CVio), i.e., $[g({\mathbf{x}})]_{+}$, and the stationarity violation (SVio). At each iteration, we solve
$\min_{{\mathbf{x}}}\big{\{}f({\mathbf{x}})+\rho_{f}\|{\mathbf{x}}-{\mathbf{x}}^{(k)}\|^{2},\,\text{{s.t.}}\,g({\mathbf{x}})+\rho_{g}\|{\mathbf{x}}-{\mathbf{x}}^{(k)}\|^{2}\leq 0\big{\}}$
@@
to a desired accuracy to obtain $\widehat{{\mathbf{x}}}^{(k)}$ and then use $\|\widehat{{\mathbf{x}}}^{(k)}-{{\mathbf{x}}}^{(k)}\|$ to measure SVio [^3] Specically, we solve the strongly convex problem by the SSG [18] to a tolerance of $10^{-2}$..
@@ 4.3 Comparisons among different methods
In Fig. 1, we compare the 3S-Econ-D method with three state-of-the-art deterministic approaches on solving Problems 1–2. It is clear that 3S-Econ-D outperforms the others in FVs and SVio values, not only converging faster but also achieving lower errors. In terms of CVio, 3S-Econ-D shows more fluctuations. Nevertheless, we observe that it remains feasible for most of the iterations. Since for the deterministic version, we can explicitly check the feasibility. Thus by keeping feasible iterates, 3S-Econ-D will significantly outperform the other three deterministic methods.
@@
To evaluate the numerical behavior of the stochastic version of our method (denoted as 3S-Econ-S), we first compare it with 3S-Econ-D. The FV and CVio values on Problems 1-2 with two datasets are shown in Fig. 2, “data passes” for the $x$-axis label is a shorthand for “data passes ($g$)” (DP ($g$)), referring to the number of times to access all the data involved in the constraint function. Despite more oscillations due to the stochasticity, 3S-Econ-S can produce comparable results as 3S-Econ-D with significantly fewer data passes.
@@
In Table 2, we show more comparison results of 3S-Econ-D and 3S-Econ-S with IPP-SSG, IPP-ConEx, SSG, and the stochastic version of SSG (denoted as SSG-S) [^4] It is worth noting that the stochastic versions of IPP-SSG and IPP-ConEx produce numerical results inferior to their deterministic counterparts when solving Problems 1–2. Thus, we only present results for their deterministic versions. This observation can be explained by the fact that, although their stochastic variants utilize the stochastic subgradient of $f$ rather than the full subgradient, both methods still require accessing all data in the constraints during each iteration to access a constraint stochastic subgradient and the constraint deterministic function value. . We terminate all the deterministic methods if SVio is less than $10^{-3}$, while for the stochastic algorithms, the threshold is set to $5\times 10^{-3}$. In addition, we impose a maximum “DP ($g$)” to 200,000 for Problem 1 and 720,000 for Problem 2. Since computing $\widehat{{\mathbf{x}}}^{(k)}$ with high precision at each iteration $k$ is time-consuming, we calculate SVio every few steps. Upon termination, in terms of “DP ($g$)” that dominate “DP ($f$)” for all methods, we observe that for Problem 1, 3S-Econ-D is 5 times faster than other deterministic methods on the a9a dataset and 2 times faster on the COMPAS dataset; for Problem 2, 3S-Econ-D is 2-15 times faster than other deterministic methods on the a9a dataset, and 1.5-2 times faster on the COMPAS dataset. Furthermore, 3S-Econ-S delivers even more impressive speedups: for Problem 1, it is 219 times faster than SSG-S on the a9a dataset and 108 times faster on the COMPAS dataset; for Problem 2, it is 466 times faster than SSG-S on the a9a dataset and 26 times faster on the COMPAS dataset. Additionally, in most cases, our methods demonstrate superior performance in terms of both FV and SVio when compared to other algorithms.
@@
Table 2: Comparisons among all the methods.
5 Conclusions
We have presented a novel single-loop first-order method for solving nonconvex nonsmooth stochastic optimization with expectation constraints. Built on an exact penalty formulation that applies a smoothed penalty function on the inequality expectation constraint, our method incorporates the SPIDER-type estimation of the constraint function into the stochastic subgradient method. Through establishing the equivalence between the near-stationary solution of the penalty problem and the near-KKT solution of the original problem, we achieve an optimal $O(\epsilon^{-4})$ iteration complexity result to produce an $(\epsilon,\epsilon)$-KKT point. In terms of sample complexity, our result is lower than a few state-of-the-art results by a factor of $\epsilon^{-2}$ in either constraint sample subgradient or constraint function values, or by a factor of $\sqrt{N}$ when the constraint is the average of $N$ sample functions. Moreover, our method delivers significantly superior numerical performance over state-of-the-art methods on solving fairness-constrained problems.
@@ Acknowledgement
We would like to express our sincere gratitude to Dr. Yankun Huang and Prof. Qihang Lin for providing their code in [18].
Appendix A Comparisons among our CQs and others
As mentioned in Remark 3.13, the strong feasibility condition is stronger than the (uniform) Slater-type CQ stated in 5. In the following lemmas, we demonstrate that 2 and 3 can be implied by the uniform Slate-type CQ under mild conditions.
Lemma A.1.
Under 1 and 5, if $0\leq g({\mathbf{x}})\leq B_{g}$, then $\min_{\boldsymbol{{\boldsymbol{\zeta}}_{g}}\in\partial g(\mathbf{x})}\left\|\boldsymbol{{\boldsymbol{\zeta}}_{g}}\right\|\geq\theta,$ where $\theta=\sqrt{2B(\overline{\rho}-\rho_{g})}$.
Proof A.2.
@@ Appendix A Comparisons among our CQs and others
Becasue $g({\mathbf{x}})\leq B_{g}$ and 5, There exists a point $\mathbf{y}$ such that $g(\mathbf{y})+\frac{\overline{\rho}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}\leq-B$. Since $g(\mathbf{z})+\frac{\overline{\rho}}{2}\|\mathbf{z}-\mathbf{x}\|^{2}$ is $(\overline{\rho}-\rho_{g})$-strongly convex, we obtain that for any $\boldsymbol{\zeta}_{g}\in\partial g(\mathbf{x})$, it holds
@@
(48) $g(\mathbf{x})+\frac{\overline{\rho}}{2}\|\mathbf{x}-\mathbf{x}\|^{2}+\left\langle\boldsymbol{\zeta}_{g},\mathbf{y}-\mathbf{x}\right\rangle+\frac{\overline{\rho}-\rho_{g}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}\leq g(\mathbf{y})+\frac{\overline{\rho}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}\leq-B.$
Applying Young’s inequality to Eq. 48 and noticing $g({\mathbf{x}})\geq 0$, we derive
@@
$-\frac{\left\|\boldsymbol{\zeta}_{g}\right\|^{2}}{2(\overline{\rho}-\rho_{g})}\leq\left\langle\boldsymbol{\zeta}_{g},\mathbf{y}-\mathbf{x}\right\rangle+\frac{\overline{\rho}-\rho_{g}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}\leq-B-g({\mathbf{x}})\leq-B.$
This indicates $\|\boldsymbol{\zeta}_{g}\|\geq\sqrt{2B(\overline{\rho}-\rho_{g})}$ and completes the proof.
Lemma A.3.
Suppose that 1 and 5 hold and $g$ is convex. Then for any ${\mathbf{x}}$ satisfying $g({\mathbf{x}})\leq B_{g}$, it holds that $\theta\textnormal{dist}({\mathbf{x}},\Omega)\leq g_{+}({\mathbf{x}})$, where $\theta$ is given in Lemma A.1.
Proof A.4.
Define $\mathbf{x}^{+}=\operatorname{Proj}_{\Omega}(\mathbf{x})\in\underset{{\mathbf{y}}:g(\mathbf{y})\leq 0}{\arg\min}\frac{1}{2}\|\mathbf{y}-\mathbf{x}\|^{2}.$ If $\mathbf{x}\in\Omega$, the result holds trivially. Therefore, we assume $\mathbf{x}\not\in\Omega$, which implies $g\left(\mathbf{x}^{+}\right)=0$. Under the Slater-type CQ, there exist $\lambda^{+}\geq 0$ and $\boldsymbol{\zeta}_{g}\in\partial g\left(\mathbf{x}^{+}\right)$ such that $\mathbf{x}^{+}-\mathbf{x}+\lambda^{+}\boldsymbol{\zeta}_{g}=\mathbf{0}.$ We assert that $\lambda^{+}>0$, because otherwise, the above inclusion would imply $\mathbf{x}=\mathbf{x}^{+}$, which contradicts the assumption that $\mathbf{x}\not\in\Omega$. By the convexity of $g$, we have
(49) $\lambda^{+}\left(g(\mathbf{x})-g\left(\mathbf{x}^{+}\right)\right)\geq\lambda^{+}\left\langle\boldsymbol{\zeta}_{g},\mathbf{x}-\mathbf{x}^{+}\right\rangle=\left\langle\lambda^{+}\boldsymbol{\zeta}_{g},\mathbf{x}-\mathbf{x}^{+}\right\rangle=\left\|\mathbf{x}-\mathbf{x}^{+}\right\|^{2}.$
Dividing both sides of the inequalities by $\lambda^{+}$ in Eq. 49 yields
$g(\mathbf{x})-g\left(\mathbf{x}^{+}\right)\geq\frac{\left\|\mathbf{x}-\mathbf{x}^{+}\right\|^{2}}{\lambda^{+}}=\|\boldsymbol{\zeta}_{g}\|\|\mathbf{x}-\mathbf{x}^{+}\|\geq\sqrt{2B(\overline{\rho}-\rho_{g})}\left\|\mathbf{x}-\mathbf{x}^{+}\right\|,$
where the last inequality is by Lemma A.1. Noting that $g(\mathbf{x})=g_{+}(\mathbf{x})$ for $\mathbf{x}\in\Omega$ and $g\left(\mathbf{x}^{+}\right)=0$, we obtain $\theta\textnormal{dist}({\mathbf{x}},\Omega)\leq g_{+}({\mathbf{x}})$, and complete the proof.
Finally, we remove the convexity assumption from the last lemma and present the following result.
Lemma A.5.
Suppose that 1 and 5 hold, and $\rho_{g}>0$. Then for any ${\mathbf{x}}$ satisfying $\textnormal{dist}({\mathbf{x}},\Omega)\leq\min\{\frac{2\gamma}{\rho_{g}},\frac{B_{g}}{l_{g}}\}$, it holds that $\gamma\textnormal{dist}({\mathbf{x}},\Omega)\leq g_{+}({\mathbf{x}})$, where $\gamma=\sqrt{\frac{B(\overline{\rho}-\rho_{g})}{2}}$.
Proof A.6.
Define $\mathbf{x}^{+}=\operatorname{Proj}_{\Omega}(\mathbf{x})\in\underset{{\mathbf{y}}:g(\mathbf{y})\leq 0}{\arg\min}\frac{1}{2}\|\mathbf{y}-\mathbf{x}\|^{2}.$ If $\mathbf{x}\in\Omega$, the result holds trivially. Therefore, we assume $\mathbf{x}\not\in\Omega$, which implies $g\left(\mathbf{x}^{+}\right)=0$. It is Straightforward that $\mathbf{x}^{+}$ is also the optimal solution of the convex problem $\underset{{\mathbf{y}}:g(\mathbf{y})+\frac{\overline{\rho}}{2}\|{\mathbf{y}}-\mathbf{x}^{+}\|^{2}\leq 0}{\min}\frac{1}{2}\|\mathbf{y}-\mathbf{x}\|^{2}$. By 5, the Slater’s condition holds for the above problem. Hence, there exist $\lambda^{+}\geq 0$ and $\boldsymbol{\zeta}_{g}\in\partial g\left(\mathbf{x}^{+}\right)$ such that $\mathbf{x}^{+}-\mathbf{x}+\lambda^{+}\boldsymbol{\zeta}_{g}=\mathbf{0}.$ We assert that $\lambda^{+}>0$, as otherwise, the above inclusion would imply $\mathbf{x}=\mathbf{x}^{+}$, which contradicts to $\mathbf{x}\not\in\Omega$. By the $\rho_{g}$-weak convexity of $g$, we have

### v2

@@ 4.1 Tested problems and Datasets
(42) $\min_{{\mathbf{x}}:\|{\mathbf{x}}\|_{\infty}\leq D}\Phi({\mathbf{x}})+\lambda\|\varphi({\mathbf{x}})\|_{1},\,\,\text{{s.t.}}\,\,\Psi_{0}({\mathbf{x}})\leq\kappa_{2},$
where $\varphi:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$, $\Psi_{0}:\mathbb{R}^{d}\mapsto\mathbb{R}$, $\Phi$ is the same as that in Eq. 5, $D>0$, $\lambda>0$, $\kappa_{2}>0$ is a slackness parameter. The functions $\Psi_{0}$ and $\varphi$ are defined as follows:
@@
$\small{[\varphi({\mathbf{x}})]_{i}=\begin{cases}2\left|x_{i}\right|,&0\leq\left|x_{i}\right|\leq 1,\\ -x_{i}^{2}+4\left|x_{i}\right|+1,&1<\left|x_{i}\right|\leq 2,\\ 3,&\left|x_{i}\right|>2,\end{cases}\quad\text{ for }i=1,2,\ldots,d},$
where $\varphi$ promotes sparsity in $\mathbf{x}$}. Here, $\Psi_{0}$ is a continuous relaxation of the demographic parity $\left|\frac{1}{n_{\mathrm{p}}}\sum_{i=1}^{n_{\mathrm{p}}}\mathbb{I}\left(\mathbf{x}^{\top}\mathbf{a}_{i}^{\mathrm{p}}\geq 0\right)-\frac{1}{n_{\mathrm{u}}}\sum_{i=1}^{n_{\mathrm{u}}}\mathbb{I}\left(\mathbf{x}^{\top}\mathbf{a}_{i}^{\mathrm{u}}\geq 0\right)\right|$. For this problem, we discuss in Appendix B that its constraint satisfies 2, and from [27], we obtain $\rho_{f}=\rho_{g}=\max\big{\{}2\lambda,\frac{1}{n_{\mathrm{p}}}\sum_{i=1}^{4n_{\mathrm{p}}}\|{\mathbf{a}}_{i}^{\mathrm{p}}\|^{2}+\frac{1}{n_{\mathrm{u}}}\sum_{i=1}^{4n_{\mathrm{u}}}\|{\mathbf{a}}_{i}^{\mathrm{u}}\|^{2}\big{\}}$, and the Lipschitz modulus $l_{g}=\frac{1}{n_{\mathrm{p}}}\sum_{i=1}^{n_{\mathrm{p}}}\|{\mathbf{a}}_{i}^{\mathrm{p}}\|+\frac{1}{n_{\mathrm{u}}}\sum_{i=1}^{4n_{\mathrm{u}}}\|{\mathbf{a}}_{i}^{\mathrm{u}}\|$, $l_{f}=5\lambda+\frac{1}{n}\sum_{i=1}^{n}\|{\mathbf{a}}_{i}\|$. In the experiments, we set $D=5$, $\lambda=0.02$, $\kappa_{2}=0.02$, and ${\mathbf{x}}^{(0)}=\mathbf{0}$ for all the compared methods.
Problem 3. The third tested problem is a multi-class Neyman-Pearson classification problem formulated in Eq. 3. We set $D=0.3$, $\kappa=4.5$, and let $\phi$ be the sigmoid function $\sigma$. Similar to those constants of Problem 2, we have $l_{f}=l_{g}=\max_{i\in[M]}\frac{1}{|{\mathcal{D}}_{i}|}\sum_{{\mathbf{a}}_{i}\in{\mathcal{D}}_{i}}\|{\mathbf{a}}_{i}\|$, and $\rho_{f}=\rho_{g}=\max_{i\in[M]}\frac{1}{|{\mathcal{D}}_{i}|}\sum_{{\mathbf{a}}_{i}\in{\mathcal{D}}_{i}}\|{\mathbf{a}}_{i}\|^{2}$. As discussed in Appendix B, the constraint of the tested problem also satisfies 2.
Datasets. The experiments of Problems 1–2 are conducted on two datasets: a9a [30] and COMPAS [5]. The a9a dataset contains 48,842 data points with 123 features and is used to predict income levels; gender is the protected attribute, and fairness is assessed between males and females. The COMPAS dataset includes 6,172 data points and 16 features, focused on predicting recidivism risk, with fairness evaluated between Caucasian and non-Caucasian groups. We split each dataset into two subsets with a ratio of 2:1. The larger set is used as $\mathcal{D}=\left\{\left(\mathbf{a}_{i},b_{i}\right)\right\}_{i=1}^{n}$, and the smaller subgroup is further partitioned into two subsets based on a binary protected attribute: $\mathcal{D}_{\mathrm{p}}=\left\{\left(\mathbf{a}^{\mathrm{p}}_{i},b^{\mathrm{p}}_{i}\right)\right\}_{i=1}^{n_{\mathrm{p}}}$ and $\mathcal{D}_{\mathrm{u}}=\left\{\left(\mathbf{a}^{\mathrm{u}}_{i},b^{\mathrm{u}}_{i}\right)\right\}_{i=1}^{n_{\mathrm{u}}}$. The experiments for Problem 3 are conducted using the MNIST dataset [34], which contains 60,000 grayscale images of handwritten digits (0–9), each image having $\times28$ pixels and an associated ground-truth label. We set $M=10$ and define each subset ${\mathcal{D}}_{i}$ as the collection of images corresponding to the image labeled by digit $\mathrm{mod}(i,10)$, for $i\in[M]$.
@@ 4.2 Implementation details
For our method, we use $\beta=10$ and $\nu=10^{-5}$ as the default setting. We implement both deterministic and stochastic versions of our method (denoted as 3S-Econ-D and 3S-Econ-S, respectively). For 3S-Econ-D, we set $q=1$ and $S_{1}=S_{2}=n$ (i.e., taking the whole data) and select $\alpha_{k}=10^{-2}$ as our default setting. For 3S-Econ-S, we set $S_{1}=n$ and $S_{2}=q=\lceil\sqrt{n}\rceil$ and select $\alpha_{k}=\frac{1}{100\lceil\sqrt{k/q}\rceil}$ as our default setting [^3] To determine the default settings of the stepsize, we test 3S-Econ-D and 3S-Econ-S on Problem 1 using the a9a dataset, with varying choices of $\alpha_{k}$. Specifically, for all $k\geq 0$, we choose $\alpha_{k}$ from $\{0.001,0.01,\frac{1}{100\lceil\sqrt{k/q}\rceil},\frac{1}{1000\lceil\sqrt{k/q}\rceil}\}$. Through our experiments, we observe that 3S-Econ-D achieves the best performance with $\alpha_{k}=0.01,\forall\,k$. Similarly, for 3S-Econ-S, $\alpha_{k}=\frac{1}{100\lceil\sqrt{k/q}\rceil},\forall\,k,$ yields the best performance. Based on these observations, we fix these parameters as the default settings and apply them across all datasets and tested problems. .
@@
We adopt the tuning techniques in [10, 46] for the two distinct IPP methods and the techniques in [27] for the SSG. A strongly convex constrained subproblem is solved approximately in each outer iteration of the IPPs. For this subproblem, we employ the SSG from [27] and the ConEx in [10] as inner solvers. We refer to the two implemented IPPs as IPP-SSG and IPP-ConEx, respectively. The stepsize $\alpha_{k}$ for IPP-SSG is selected from $\{2\times 10^{-4},5\times 10^{-4},10^{-3},5\times 10^{-3}\}$. Following the notation in [46], we set $\theta_{t}=\frac{t}{t+1},\eta_{t}=c_{1}(t+1)$, and $\tau_{t}=\frac{c_{2}}{t+1}$ for IPP-ConEx, where $c_{1}$ is chosen from $\{20,50,100,200\}$ and $c_{2}$ from $\{0.002,0.005,0.01,0.02\}$. The optimal parameter set for these methods is selected based on the smallest objective value after 5000 iterations.
@@
We report the objective function value (FV), i.e., $f({\mathbf{x}})$, the constraint violation (CVio), i.e., $\|[{\mathbf{g}}({\mathbf{x}})]_{+}\|_{1}$, and the stationarity violation (SVio). At each iteration, we solve
$\min_{{\mathbf{x}}}\big{\{}f({\mathbf{x}})+\rho_{f}\|{\mathbf{x}}-{\mathbf{x}}^{(k)}\|^{2},\,\text{{s.t.}}\,g_{i}({\mathbf{x}})+\rho_{g}\|{\mathbf{x}}-{\mathbf{x}}^{(k)}\|^{2}\leq 0,\,i\in[m]\big{\}}$
@@
to a desired accuracy to obtain $\widehat{{\mathbf{x}}}^{(k)}$ and then use $\|\widehat{{\mathbf{x}}}^{(k)}-{{\mathbf{x}}}^{(k)}\|$ to measure SVio [^4] Specifically, we solve the strongly convex problem by the SSG [27] to a tolerance of $10^{-2}$..
4.3 Performance of 3S-Econ-D with different $\beta$ and $\nu$
In this subsection, we investigate the numerical performance of 3S-Econ-D with different values of $\beta$ and $\nu$ on solving Problem 1 by using the a9a dataset, respectively. We vary the penalty parameter $\beta\in\{1,10,100,1000\}$ with fixed $\nu=10^{-5}$, and subsequently vary the smoothing parameter $\nu\in\{10^{-5},10^{-4},10^{-3},10^{-2}\}$ with fixed $\beta=10$. The results are depicted in Fig. 1. From the figure, we observe that when $\beta=1$, the CVio remains relatively large, indicating inadequate penalization, while excessively large $\beta$ slows down convergence. Among the varying values, $\beta=10$ yields the most balanced trade-off between convergence speed and constraint satisfaction. In addition, we observe that the CVio closely aligns with $\nu$, consistent with our theoretical results. On all subsequent experiments, we select $\beta=10$ and $\nu=10^{-5}$ as our default settings.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 1: Performance of 3S-Econ-D with different $\beta$ and $\nu$ on solving Problem 1 by using the a9a dataset. Top row: different values of $\beta$ and fixed $\nu=10^{-5}$; Bottom row: different values of $\nu$ and fixed $\beta=10$.
@@ 4.4 Comparisons among different methods
In Fig. 2, we compare our 3S-Econ-D method with three state-of-the-art deterministic approaches on solving Problems 1–2 over iterations [^5] Here, for the two IPP methods, the iteration count refers specifically to the inner-loop iterations.. It is clear that 3S-Econ-D outperforms the others in FVs and SVio values, not only converging faster but also achieving lower errors. In terms of CVio, 3S-Econ-D shows more fluctuations. Nevertheless, we observe that it remains feasible for most of the iterations. Since for the deterministic version, we can explicitly check the feasibility. By keeping feasible iterates, 3S-Econ-D will significantly outperform the other three deterministic methods.
@@
To evaluate the numerical behavior of the stochastic version of our method (denoted as 3S-Econ-S), we first compare it with 3S-Econ-D. The FV and CVio values on Problems 1-2 with two datasets are shown in Fig. 3, “data passes” for the $x$-axis label is a shorthand for “data passes ($\mathbf{g}$})” (DP ($\mathbf{g}$})), referring to the number of times to access all the data involved in the constraint function. Despite more oscillations due to the stochasticity, 3S-Econ-S can produce comparable results to 3S-Econ-D with significantly fewer data passes.
@@
In Table 2, we show more comparison results of 3S-Econ-D and 3S-Econ-S with IPP-SSG, IPP-ConEx, SSG, and the stochastic version of SSG (denoted as SSG-S) [^6] It is worth noting that the stochastic versions of IPP-SSG and IPP-ConEx produce numerical results inferior to their deterministic counterparts when solving Problems 1–2. Thus, we only present results for their deterministic versions. This observation can be explained by the fact that, although their stochastic variants utilize the stochastic subgradient of $f$ rather than the full subgradient, both methods still require accessing all data in the constraints during each iteration to access a constraint stochastic subgradient and the constraint deterministic function value. . We terminate all the deterministic methods if SVio is less than $10^{-3}$, while for the stochastic algorithms, the threshold is set to $5\times 10^{-3}$. In addition, we impose a maximum “DP ($\mathbf{g}$})” of 200,000 for Problem 1 and 720,000 for Problem 2. Since computing $\widehat{{\mathbf{x}}}^{(k)}$ with high precision at each iteration $k$ is time-consuming, we calculate SVio every few steps. Upon termination, our proposed methods demonstrate superior performance in terms of both FV and SVio compared to other tested algorithms in most cases. In terms of “DP ($\mathbf{g}$})”, which consistently dominates “DP (ff)” for all methods, we observe that for Problem 1, 3S-Econ-D achieves speedups ranging from 1.5 to 5 times compared to other deterministic algorithms, while 3S-Econ-S delivers even more impressive speedups, approximately 108 to 466 times faster. When considering the metric “(CPU) Time (s)”, both 3S-Econ-D and 3S-Econ-S remain significantly faster than their counterparts. Although the acceleration factor of 3S-Econ-S in terms of runtime is less pronounced compared to that measured by “DP ($\mathbf{g}$})”, this discrepancy primarily arises from implementation-related overhead inherent in MATLAB, such as matrix-vector multiplications.
@@
Table 2: Comparisons among different methods on solving Problems 1 and 2.
In Fig. 4, we solve Problem 3 on the MNIST dataset and present the values of FV, CVio, and SVio versus CPU time for SSG, IPP-ConEx, IPP-SSG, and our proposed methods. Overall, our methods reach lower FV and SVio values more quickly compared to other methods, while achieving feasibility at most iterates. Specifically, 3S-Econ-S attains an SVio of $1.29\times 10^{-3}$ at 21.9 seconds, and 3S-Econ-D reaches an SVio of $1.53\times 10^{-3}$ at 172 seconds, significantly outperforming SSG, which achieves its lowest SVio of $1.07\times 10^{-2}$ at about 225 seconds. Additionally, IPP-ConEx remains stagnant at the zero solution, likely due to the zero point being a stationary solution of its proximal subproblem.
Figure 4: Comparisons of different methods on Problem 3 by using MNIST dataset.
5 Conclusions and Discussions
We have presented a novel FOM for solving nonconvex nonsmooth stochastic optimization with expectation constraints. Built on an exact penalty formulation that applies a smoothed penalty function on the inequality expectation constraint, our method incorporates the SPIDER-type estimation of the constraint function into the stochastic subgradient method. Through establishing the equivalence between the near-stationary solution of the penalty problem and the near KKT solution of the original problem, we achieve an optimal $O(\epsilon^{-4})$ iteration complexity result to produce an $(\epsilon,\epsilon)$-KKT point. In terms of sample complexity, our result is lower than a few state-of-the-art results by a factor of $\epsilon^{-2}$ in either constraint sample subgradient or constraint function values, or by a factor of $\sqrt{N}$ when the constraint is the average of $N$ sample functions. Moreover, our method delivers significantly superior numerical performance over state-of-the-art methods on solving fairness-constrained problems and Neyman-Pearson classification problem.
[…]
@@ Acknowledgement
We would like to express our sincere gratitude to Dr. Yankun Huang and Prof. Qihang Lin for providing their code in [27]. We also thank the three anonymous referees and the associate editor for their constructive comments.
Appendix A Proof of Lemma 2.4
Proof A.1.
@@
Let $\mathcal{J}_{3}=\{i:g_{i}({\mathbf{x}})\geq 0\}$. Then by the condition on ${\mathcal{J}}_{1}$, it holds $\mathcal{J}_{1}\cap\mathcal{J}_{3}\neq\emptyset$, and the set difference $\mathcal{J}_{1}\cap\mathcal{J}_{3}^{c}\subset\{i:g_{i}({\mathbf{x}})<0\}$. Since $0<g_{+}({\mathbf{x}})\leq B_{g}$, then by 2, there exists $\mathbf{y}\in\mathrm{relint}(\mathcal{X})$ such that $g_{i}(\mathbf{y})+\frac{\overline{\rho}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}\leq-B$ for all $i\in[m]$. Now, for any $\boldsymbol{\zeta}_{g_{i}}\in\partial[g_{i}({\mathbf{x}})]_{+},{i\in\mathcal{J}_{1}}$, any $\boldsymbol{\zeta}_{g_{i}}\in\partial g_{i}({\mathbf{x}}),{i\in\mathcal{J}_{2}}$, and any ${\mathbf{u}}\in\mathcal{N}_{\mathcal{X}}({\mathbf{x}})$, we have
@@
$\sum_{i\in\mathcal{J}_{1}}[g_{i}(\mathbf{x})]_{+}+\sum_{i\in\mathcal{J}_{2}}\lambda_{i}g_{i}(\mathbf{x})+\left\langle\sum_{i\in J_{1}}{\boldsymbol{\zeta}}_{g_{i}}+\sum_{i\in{\mathcal{J}}_{2}}\lambda_{i}{\boldsymbol{\zeta}}_{g_{i}}+{\mathbf{u}},\mathbf{y}-\mathbf{x}\right\rangle+\frac{\overline{\rho}-\rho_{g}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}$
@@
$=$ $\sum_{i\in\mathcal{J}_{1}\cap{\mathcal{J}}_{3}}\left(\textstyle g_{i}(\mathbf{x})+\frac{\overline{\rho}}{2}\|\mathbf{x}-\mathbf{x}\|^{2}\right)+\textstyle\left\langle\sum_{i\in{\mathcal{J}}_{1}\cap{\mathcal{J}}_{3}}{\boldsymbol{\zeta}}_{g_{i}}+{\mathbf{u}},\mathbf{y}-\mathbf{x}\right\rangle+\frac{\overline{\rho}-\rho_{g}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}$